# Data-driven $h^{2}$ model reduction for linear discrete-time systems

### v2

@@ IV-B Construction of Gradient-like Matrices from Measurement Data
In this subsection, we show that the gradient-like matrices corresponding to Proposition 2 can be derived from the data. To obtain the gradient-like matrices, we solve the discrete-time Sylvester equations (8) and (9) in a data-driven manner, following [37][38]. Multiplying (8) and (9) by the data matrix $X$ from the left, from (18), (21), (22), and (24), we obtain
@@
Let $Z_{A,\text{data}}=Z_{A}$, $Z_{B,\text{data}}=Z_{B}$, and $U_{B,\text{data}}=U_{B}$, where $Z_{A,\text{data}}$, $Z_{B,\text{data}}$, and $U_{B,\text{data}}$ are the solutions of (21), (22), and (24), respectively. Suppose that Assumptions (A1)-(A3), and (A5) hold. Then, the solutions $R$ and $S$ to (27) and (28) exist uniquely.
@@
It is shown using the well-known unique solution existence theorem for the Sylvester equation [39]. See Appendix. $\Box$
@@
~$\tilde{\nabla}_{\hat{A}}f(\hat{A},\hat{B},\hat{C}):=2(Q\hat{A}P+(S_{\text{data}}^{\top}R_{\text{data}}-S_{B,\text{data}}^{\top}\hat{B}^{\top})(\hat{A}^{\dagger})^{\top}),$ (29)
~$\tilde{\nabla}_{\hat{B}}f(\hat{A},\hat{B},\hat{C}):=2(S_{B,\text{data}}^{\top}+Q\hat{B}),$ (30)
~$\tilde{\nabla}_{\hat{C}}f(\hat{A},\hat{B},\hat{C}):=2(\hat{C}P-R_{\text{data}}),$ (31)
@@ IV-C Data-driven $h^{2}$ Model Reduction
This subsection describes the data-driven model reduction for the case where $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ are unknown. To solve (10) numerically, we compute the gradients of Proposition 2 in a data-driven manner.
@@
Let $NL-N\geq n+m$. Suppose that Assumptions (A4)-(A6) hold. Then, (20) has the unique solution $(Z_{A,\text{data}},Z_{B,\text{data}})=(Z_{A},Z_{B})$.
Proof.
We use the condition for the existence of a unique solution to the linear equation. See Appendix. $\Box$
Lemma 2 means that under assumptions, $Z_{A}$ given by (21) and $Z_{B}$ given by (22) can be expressed using input and state data of system (1). In other words, $Z_{A}$ and $Z_{B}$ can be expressed as $Z_{A,\text{data}}$ and $Z_{B,\text{data}}$ without using the unknown matrices $A$ and $B$.
@@
Let $NL-N\geq n+m$. Suppose that Assumptions (A4)-(A6) hold. Then, (23) and (32) have the unique solutions (24) and
@@
Proof.
It is proved following the same approach as in Lemma 2. See Appendix. $\Box$
@@
Let $NL-N\geq n+m$. Suppose that Assumptions (A1)-(A6) hold. Then, the gradient-like matrices of (29)-(31) coincide with the gradients for (10) specified in Proposition 2, that is, ~$\tilde{\nabla}_{\hat{A}}f=\nabla_{\hat{A}}f$, ~$\tilde{\nabla}_{\hat{B}}f=\nabla_{\hat{B}}f$, and ~$\tilde{\nabla}_{\hat{C}}f=\nabla_{\hat{C}}f$ hold.
@@
Theorem 1 states that the gradients for (10) are computed based on the data given by (17) under Assumptions (A1)-(A6). In fact, $R_{\text{data}}$, $S_{\text{data}}$, and $S_{B,\text{data}}$ in (29), (30), and (31) can be calculated by solving (27), (28), and (32). Moreover, $(Z_{A},Z_{B})$ and $U_{B}$ in (27), (28), and (32) can be calculated by solving (20) and (23). As mentioned earlier, if Assumptions (A1)-(A6) are not satisfied, it is possible that the gradient obtained from the data differs from the true gradients as shown in Proposition 2. When Assumptions (A1)-(A6) regarding the data are satisfied and the gradients from Theorem 1 are all zero, the stationary point condition described in Section II is satisfied. This fact becomes crucial in the construction of the proposed algorithm, as discussed in Section V.
@@
Our framework assumes that the full state is directly measured, that is, $p=n$ and, without loss of generality, $C=I_{n}$. Under this setting we have $y_{i,k}=x_{i,k}$, and therefore the state sequence is exactly known and Theorem 1 provides the true gradients of (10). If only a part of the state is observed ($p<n$) or, more generally, if $C\neq I_{n}$, the state $x_{i,k}$ must be reconstructed, which inevitably introduces approximation errors. In such a partially observed setting the matrices ~$\tilde{\nabla}_{\hat{A}}f$, ~$\tilde{\nabla}_{\hat{B}}f$, and ~$\tilde{\nabla}_{\hat{C}}f$ no longer coincide with the true gradients. Future work will address this case.
@@ IV-D Gradients of (10) for the case that $A$ is unknown and $B$ is known
Let $NL-N\geq n$. Suppose that Assumptions (A1)-(A3), and (A5) hold. Then, the gradient-like matrices of (29)-(31) coincide with the gradients for (10) specified in Proposition 2, that is, ~$\tilde{\nabla}_{\hat{A}}f=\nabla_{\hat{A}}f$, ~$\tilde{\nabla}_{\hat{B}}f=\nabla_{\hat{B}}f$, and ~$\tilde{\nabla}_{\hat{C}}f=\nabla_{\hat{C}}f$ hold true.
@@
Under Assumptions (A1)-(A6), the gradients (29)-(31) with unknown $B\in\mathbb{R}^{n\times m}$ coincide with the gradients with known $B\in\mathbb{R}^{n\times m}$.
@@ V Gradient-based data-driven $h^{2}$ MOR algorithm
In this section, we present Algorithm 1 which solves the problem (10) when $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ are unknown.
We briefly describe Algorithm 1. First, initial reduced matrices $\hat{\theta}_{1}=(\hat{A}_{(1)},\hat{B}_{(1)},\hat{C}_{(1)})$ are created to satisfy Assumption (A3). In step 1 and 2 of Algorithm 1, $Z_{A,\text{data}}$, $Z_{B,\text{data}}$, and $U_{B,\text{data}}$, which are needed to solve the discrete-time Sylvester equations (27) and (28) with $(Z_{A},Z_{B},U_{B})=(Z_{A,\text{data}},Z_{B,\text{data}},U_{B,\text{data}})$, are generated. At each iteration, after computing the gradients, the reduced matrices are updated so that the objective function for (10) becomes smaller. In the while statement of Algorithm 1, the backtracking method is executed.
Here, we show that for a sufficiently small step size $\alpha$, the updated matrix $\bar{A}$ satisfies Assumption 3. Since the set of stable matrices $\mathcal{\hat{D}}_{\text{stab}}:=\{\hat{A}\in\mathbb{R}^{r\times r}|\forall i\in\{1,2,\ldots,r\},|\lambda_{i}(\hat{A})|<1\}$ is open [40], and the set of nonsingular matrices $\mathcal{\hat{D}}_{\text{nonsing}}:=\{\hat{A}\in\mathbb{R}^{r\times r}|\forall i\in\{1,2,\ldots,r\},|\lambda_{i}(\hat{A})|>0\}$ is also open due to the continuity of the determinant, their intersection $\mathcal{\hat{D}}:=\mathcal{\hat{D}}_{\text{stab}}\bigcap\mathcal{\hat{D}}_{\text{nonsing}}=\{\hat{A}\in\mathbb{R}^{r\times r}|\forall i\in\{1,2,\ldots,r\},0<|\lambda_{i}(\hat{A})|<1\}$ is open by the basic topological axiom that the finite intersection of open sets is open. Hence $\mathcal{\hat{D}}$, which denotes the set of matrices satisfying Assumption 3, is an open subset of $\mathbb{R}^{r\times r}$. Thus, there exists $\varepsilon(\hat{A}_{(\ell)})>0$ such that $\left\{\bar{A}\in\mathbb{R}^{r\times r}\,\middle|\,\|\bar{A}-\hat{A}_{(\ell)}\|_{F}<\varepsilon(\hat{A}_{(\ell)})\right\}\subset\hat{\mathcal{D}}$. Hence for the gradient $\nabla_{\hat{A}}f(\hat{A}_{(\ell)})$, we choose
$0<\alpha_{\ell}<\alpha^{\star}_{\ell}\quad\text{with}\quad\alpha^{\star}_{\ell}:=\frac{\varepsilon(\hat{A}_{(\ell)})}{\|\nabla_{\hat{A}}f(\hat{A}_{(\ell)})\|_{F}}.$
Then $\|\hat{A}_{(\ell)}-\alpha_{\ell}\nabla_{\hat{A}}f(\hat{A}_{(\ell)})-\hat{A}_{(\ell)}\|_{F}=\alpha_{\ell}\|\nabla_{\hat{A}}f(\hat{A}_{(\ell)})\|_{F}<\varepsilon(\hat{A}_{(\ell)})$ and therefore $\bar{A}:=\hat{A}_{(\ell)}-\alpha_{\ell}\nabla_{\hat{A}}f(\hat{A}_{(\ell)})\in\hat{\mathcal{D}}.$
In addition, since $c>0$ and $D>0$, there exists $\alpha$ satisfying the Armijo condition. Furthermore, from a numerical perspective, the spectrum of $\hat{A}^{-1}$ does not intersect with that of $X^{\dagger}Z_{A,\text{data}}$ and $X^{\dagger}(X_{\mathrm{p}}-U_{B,\text{data}})$. That is, $\alpha$ can be determined by the backtracking method such that the Armijo condition and Assumptions (A1)- (A3) are satisfied.
Remark 4.
Since (10) is a non-convex optimization problem, the choice of the initial point is important. As the initial point for Algorithm 1, it is possible to use ROMs obtained through the following existing data-driven model reduction methods.
  • DMDc [27][28]
  • Loewner framework [17, 22, 21, 20, 19]
  • Data-driven balanced truncation [15, 16, 14]
Using ROMs obtained from these methods as initial points for Algorithm 1 often enables the construction of high-quality ROMs in the sense of the $h^{2}$ norm. However, these data-driven methods require different types of data, and therefore, they may not be directly applicable in some cases.
The convergence of Algorithm 1 is guaranteed by the following proposition.
@@
Suppose that the sequence $\hat{\theta}_{\ell}:=(\hat{A}_{(\ell)},\hat{B}_{(\ell)},\hat{C}_{(\ell)})$ generated by Algorithm 1 is bounded. In addition, we assume that each step-size $\alpha_{\ell}$ is chosen by the backtracking method with parameters $10<c_{1},\rho<$ 1 in Algorithm 1. Then the whole sequence $\{\hat{\theta}_{\ell}\}$ converges to a stationary point of problem (10).
Proof.
The backtracking method gives $f(\hat{\theta}_{\ell+1})\leq f(\hat{\theta}_{\ell})-c_{1}\alpha_{\ell}\cdot\|\nabla f(\hat{\theta}_{\ell})\|^{2}_{F}$, therefore $\{f(\hat{\theta}_{\ell})\}$ is monotone and bounded. Summation yields $\sum_{\ell=1}^{\infty}\alpha_{\ell}\|\nabla f(\hat{\theta}_{\ell})\|_{F}^{2}<\infty$.
Since $f$ is continuously differentiable and $10<\rho<$ 1, there exists the lower bound $\alpha_{\ell}\geq\underline{\alpha}>0$[41]. Hence $\underline{\alpha}\cdot\sum_{\ell=1}^{\infty}\|\nabla f(\hat{\theta}_{\ell})\|_{F}^{2}<\infty$ and therefore $\|\nabla f(\hat{\theta}_{\ell})\|_{F}\to 0$.
Because $\|\hat{\theta}_{\ell+1}-\hat{\theta}_{\ell}\|_{F}=\alpha_{\ell}\|\nabla f(\hat{\theta}_{\ell})\|_{F}$, $\{\hat{\theta}_{\ell}\}$ is Cauchy sequence and converges. Continuity of $\nabla f$ implies that the limit point $\hat{\theta}^{\star}$ satisfies $\nabla f(\hat{\theta}^{\star})=0$, i.e. it is stationary. $\Box$
Lastly, we provide the computational complexity analysis of Algorithm 1.
Theorem 2.
We assume that the number of trials of the backtracking method and the number of iterations of Algorithm 1 are $\mathcal{O}(k)$ and $\mathcal{O}(l)$, respectively. Then, the overall computational complexity of Algorithm 1 is given by
$\mathcal{O}((NL-N)^{2}(n+m)+((NL-N)n^{2}+n^{3}+r^{3})kl).$
@@
A widely–used alternative for constructing a ROM directly from snapshot data is DMDc [28, 27]. Given the snapshot matrix $X\in\mathbb{R}^{(NL-N)\times n}$, $X_{\text{p}}\in\mathbb{R}^{(NL-N)\times n}$, $U\in\mathbb{R}^{(NL-N)\times m}$ by (17), the computational complexity by DMDc [28, 27] is $\mathcal{O}\!\bigl{(}(NL-N)^{2}(r+m)+(NL-N)nr\bigr{)}$. Although Theorem 2 indicates that the proposed method incurs a higher computational complexity than DMDc, it can generate a stable ROM with a smaller $h^{2}$ error, as demonstrated in Section VI.
Algorithm 1 Solve (10) for the case that $A$ and $B$ are unknown
0:  Measurement data $(x_{i,k}$, $u_{i,k})$, $(i=1,2,\ldots,N,k=1,2,\ldots,L)$, dimension $r$ for (2), initial step-size $\alpha$, Armijo parameter $c_{1}$, search control parameter $\rho$, tolerance $tol$, the initial real matrices $\hat{\theta}_{1}=(\hat{A}_{(1)},\hat{B}_{(1)},\hat{C}_{(1)})$ which satisfy Assumption (A3)
0:  $\hat{\theta}=(\hat{A},\hat{B},\hat{C})\in\mathbb{R}^{r\times r}\times\mathbb{R}^{r\times m}\times\mathbb{R}^{n\times r}$
@@
6:     Calculate the gradients $\nabla_{\hat{A}}f(\hat{\theta}_{\ell})$, $\nabla_{\hat{B}}f(\hat{\theta}_{\ell})$, and $\nabla_{\hat{C}}f(\hat{\theta}_{\ell})$ with $S_{B}$ obtained by (32)
7:     Define $\nabla f(\hat{\theta}_{\ell}):=-(\nabla_{\hat{A}}f(\hat{\theta}_{\ell}),\nabla_{\hat{B}}f(\hat{\theta}_{\ell}),\nabla_{\hat{C}}f(\hat{\theta}_{\ell}))$
8:     if $\|\nabla f(\hat{\theta}_{\ell})\|_{F}^{2}<tol$ then break
9:     $\alpha_{\ell}=\alpha$
10:     while 1 do
11:        $\bar{\theta}:=\hat{\theta}_{\ell}+\alpha_{\ell}d$
@@
14:        if $f(\bar{\theta})\leq f(\hat{\theta}_{\ell})-c_{1}\alpha_{\ell}\cdot\|\nabla f(\hat{\theta}_{\ell})\|_{F}^{2}$ and $\forall i,0<|\lambda_{i}(\bar{A})|<1$ then $\hat{\theta}_{\ell+1}:=\bar{\theta}$ break

### v3

@@ IV-B Construction of Gradient-like Matrices from Measurement Data
In this subsection, we show that the gradient-like matrices corresponding to Proposition 2 can be derived from the data. To obtain the gradient-like matrices, we solve the discrete-time Sylvester equations (9) and (10) in a data-driven manner, following [34][35]. Multiplying (9) and (10) by the data matrix $X$ from the left, from (19), (22), and (24), we obtain
@@
Let $Z_{A,\text{data}}=Z_{A}$, $Z_{B,\text{data}}=Z_{B}$, and $U_{B,\text{data}}=U_{B}$, where $Z_{A,\text{data}}$, $Z_{B,\text{data}}$, and $U_{B,\text{data}}$ are the solutions of (22) and (24), respectively. Suppose that (A1)-(A3), and (A5) hold. Then, the solutions $R$ and $S$ to (27) and (28) exist uniquely.
@@
It is shown using the unique solution existence theorem for the Sylvester equation [36]. See Appendix. $\Box$
@@
$\tilde{\nabla}_{\hat{A}}f(\hat{A},\hat{B},\hat{C}):=2(Q\hat{A}P+(S_{\text{data}}^{\top}R_{\text{data}}-S_{B,\text{data}}^{\top}\hat{B}^{\top})(\hat{A}^{\dagger})^{\top}),$ (29)
[…]
$\tilde{\nabla}_{\hat{C}}f(\hat{A},\hat{B},\hat{C}):=2(\hat{C}P-R_{\text{data}}),$ (31)
@@ IV-C Data-driven $h^{2}$ Model Reduction
This subsection describes the data-driven model reduction for the case where $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ are unknown. To solve (11) numerically, we compute the gradients of Proposition 2 in a data-driven manner. We first introduce two lemmas, whose proofs are in the appendix.
@@
Let $NL-N\geq n+m$. Suppose that (A4)-(A6) hold. Then, (21) has the unique solution $(Z_{A,\text{data}},Z_{B,\text{data}})=(Z_{A},Z_{B})$.
Lemma 2 means that under assumptions, $Z_{A}$ and $Z_{B}$ given by (22) can be expressed using input and state data of system (1). In other words, $Z_{A}$ and $Z_{B}$ can be expressed as $Z_{A,\text{data}}$ and $Z_{B,\text{data}}$ without using the unknown matrices $A$ and $B$.
@@
Let $NL-N\geq n+m$. Suppose that (A4)-(A6) hold. Then, (23) and (32) have the unique solutions (24) and
@@
Let $NL-N\geq n+m$. Suppose that (A1)-(A6) hold. Then, the gradient-like matrices of (29)-(31) coincide with the gradients for (11) specified in Proposition 2, that is, $\tilde{\nabla}_{\hat{A}}f=\nabla_{\hat{A}}f$, $\tilde{\nabla}_{\hat{B}}f=\nabla_{\hat{B}}f$, and $\tilde{\nabla}_{\hat{C}}f=\nabla_{\hat{C}}f$ hold.
@@
Theorem 1 states that the gradients for (11) are computed based on the data given by (18) under (A1)-(A6). In fact, $R_{\text{data}}$, $S_{\text{data}}$, and $S_{B,\text{data}}$ in (29), (30), and (31) can be calculated by solving (27), (28), and (32). Moreover, $(Z_{A},Z_{B})$ and $U_{B}$ in (27), (28), and (32) can be calculated by solving (21) and (23). As mentioned earlier, if (A1)-(A6) are not satisfied, it is possible that the gradients obtained from the data differ from the true gradients as shown in Proposition 2. When (A1)-(A6) regarding the data are satisfied and the gradients from Theorem 1 are all zero, the stationary point condition described in Section II is satisfied. This fact becomes crucial in the construction of the proposed algorithm, as discussed in Section V.
@@
Our framework assumes that the full state is directly measured, that is, $p=n$ and, without loss of generality, $C=I_{n}$. Under this setting we have $y_{i,k}=x_{i,k}$, and therefore the state sequence is exactly known and Theorem 1 provides the true gradients of (11). If only a part of the state is observed ($p<n$) or, more generally, if $C\neq I_{n}$, the state $x_{i,k}$ must be reconstructed, which inevitably introduces approximation errors. In such a partially observed setting the matrices $\tilde{\nabla}_{\hat{A}}f$, $\tilde{\nabla}_{\hat{B}}f$, and $\tilde{\nabla}_{\hat{C}}f$ no longer coincide with the true gradients. Future work will address this case.
@@ IV-D Gradients of (11) for the case that $A$ is unknown and $B$ is known
Let $NL-N\geq n$. Suppose that (A1)-(A3), and (A5) hold. Then, the gradient-like matrices of (29)-(31) coincide with the gradients for (11) specified in Proposition 2, that is, $\tilde{\nabla}_{\hat{A}}f=\nabla_{\hat{A}}f$, $\tilde{\nabla}_{\hat{B}}f=\nabla_{\hat{B}}f$, and $\tilde{\nabla}_{\hat{C}}f=\nabla_{\hat{C}}f$ hold true.
@@
Under (A1)-(A6), the gradients (29)-(31) with unknown $B\in\mathbb{R}^{n\times m}$ coincide with the gradients with known $B\in\mathbb{R}^{n\times m}$.
@@ V Gradient-based data-driven $h^{2}$ MOR algorithm
In this section, we present the Algorithm 1 which solves the problem (11) when $A\in\mathbb{R}^{n\times n}$ and $B\in\mathbb{R}^{n\times m}$ are unknown. Consider the set $\mathcal{E}:=\{\hat{\theta}=(\hat{A},\hat{B},\hat{C})\mid 0<|\lambda_{i}(\hat{A})|<1,\ i=1,\dots,r\}$ which satisfy (A3). Since $\mathcal{E}$ is open (proved in Proposition 3 below), if the gradients from Theorem 1 are all zero at a point in $\mathcal{E}$, that point is a stationary point.
[…]
@@
Suppose $\hat{\theta}_{\ell}\in\mathcal{E}$. Then there exists $\alpha^{\star}_{\ell}>0$ such that for all $\alpha\in(0,\alpha^{\star}_{\ell})$,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\theta}:=\hat{\theta}_{\ell}-\alpha\,\tilde{\nabla}f(\hat{\theta}_{\ell})\in\mathcal{E}.}$
@@
The step-size $\alpha$ can be determined by the backtracking method such that the Armijo rule $f(\hat{\theta}_{\ell+1})\leq f(\hat{\theta}_{\ell})-c_{1}\alpha_{\ell}\>\|\tilde{\nabla}f(\hat{\theta}_{\ell})\|^{2}$ and (A1)-(A3) are satisfied. From a numerical perspective, by continuity of eigenvalues, the spectrum of $\hat{A}^{-1}$ does not intersect with that of $X^{\dagger}Z_{A,\text{data}}$ nor with that of $X^{\dagger}(X_{\mathrm{p}}-U_{B,\text{data}})$, thus (A1) and (A2) are readily satisfied. Furthermore, from Theorem 1, Proposition 3, and the smoothness of $f$, there exists $\alpha>0$ satisfying the Armijo rule and preserving feasibility.
The convergence of Algorithm 1 is guaranteed by the following theorem. Note that (A1)-(A3) are satisfied in each iteration.
Theorem 2.
[…]
Proof.
It follows from [37, Thm. 3.2]. See Appendix. $\Box$
Note that while the existence of a stationary point for (11) is non-trivial, it is guaranteed when the assumptions of Theorem 2 are satisfied. Next, we provide the computational complexity analysis of Algorithm 1.
Theorem 3.
Assume that in Algorithm 1 the backtracking terminates in at most $J$ trials per iteration, and the algorithm performs $\ell_{\text{max}}$ iterations in total. Then, the overall computational complexity of Algorithm 1 is given by
$\mathcal{O}((NL-N)^{2}(n+m)+((NL-N)n^{2}+n^{3}+r^{3})\>J\>\ell_{\text{max}}).$
Proof.
See Appendix. $\Box$
A widely–used alternative for constructing a ROM directly from snapshot data is DMDc [22, 21]. Given the snapshot matrix $X$, $X_{\text{p}}$, $U$ by (18), the computational complexity by DMDc [22, 21] is $\mathcal{O}\bigl((NL-N)^{2}(r+m)+(NL-N)nr\bigr)$. Although Theorem 3 indicates that the proposed method incurs a higher computational complexity than DMDc, it can generate a stable ROM with a smaller $h^{2}$ error, as described in Section VI.
Remark 3.
Since (11) is a nonconvex optimization problem, the choice of the initial point for Algorithm 1 is important. As the initial point, it is possible to use ROMs obtained via existing data-driven MOR methods, such as DMDc [21][22], Loewner framework [17, 18], and data-driven BT [15, 16, 14]. Such initialization often produces high-quality ROMs in the $h^{2}$-norm. However, these methods require different types of data, and therefore, they may not be directly applicable in some cases.
Remark 4.
[…]
Algorithm 1 Solve (11) for the case that AA and $B$ are unknown
0: Measurement data $(x_{i,k}$, $u_{i,k})$, $(i=1,2,\ldots,N,k=1,2,\ldots,L)$, dimension $r$ for (2), initial step-size $\alpha_{\text{init}}$, Armijo parameter $c_{1}$, search control parameter $\rho$, tolerance toltol, the initial real matrices $\hat{\theta}_{1}$ which satisfy (A3)
[…]
@@
[…]
7:  Define $)\tilde{\nabla}f(\hat{\theta}_{\ell}):=(\tilde{\nabla}_{\hat{A}}f(\hat{\theta}_{\ell}),\tilde{\nabla}_{\hat{B}}f(\hat{\theta}_{\ell}),\tilde{\nabla}_{\hat{C}}f(\hat{\theta}_{\ell}$))
8:  if $\|\tilde{\nabla}f(\hat{\theta}_{\ell})\|_{F}^{2}<tol$ then break
9:  $\alpha_{\ell}=\alpha_{\text{init}}$
10:  while true do
[…]
@@
14:   if $f(\bar{\theta})\leq f(\hat{\theta}_{\ell})-c_{1}\alpha_{\ell}\>\|\tilde{\nabla}f(\hat{\theta}_{\ell})\|_{F}^{2}$ and $\forall i\>,\>0<|\lambda_{i}(\bar{A})|<1$ then $\hat{\theta}_{\ell+1}:=\bar{\theta}$ break